40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2610     Entry time: Fri Jan 15 09:54:18 2021
Author: aaron 
Type: HowTo 
Category: PSOMA 
Subject: mode matching 

[aaron, rana]

aaron enter about Fri Jan 15 09:54:39 2021. rana about

Aaron wiped down keyboard, doorknob, light switch, other high contact surfaces. Rana pointed out several improvements we should make to the optics and electronics, and we tried to collimate the beam coming out of the fiberport.

We noticed the fiber enters the port at an angle, and tried to correct for this to no avail. Turns out, the reason for the angle is to compensate for the angle on the end of FC/APC connectors. You can see an example in the diagram for Thorlabs' fixed focus FC/APC collimators (as opposed to the FC/PC collimators, where the end is perpendicular to the beam and the fiber enters at normal incidence).

We also were concerned that we were seeing only 1.3 mW out of the fiberport -- but we were measuring the W laser path, so this measurement is consistent with the power output from that laser. The E path has 3.9 mW from the fiberport, which I think is consistent with expected losses from the 10% pickoff and couplers.

[aaron]

accounting for observed power

I'm tracing our optical path and accounting for losses. We should at some point rename the E and W lasers, since they are no longer pumping parallel cavities. Pump and signal would be appropriate; cardinal directions are ambiguous, since the relative orientation of the beams changes before the MZ.

Which laser Power after diode (mW) Before isolator After isolator Before EOAM After EOAM Before EOPM After EOPM Before 10% pickoff After 10% pickoff After fiberport
East 21.5 21.5 18.3 18.2 8.4 8.3 4.6 4.6 3.9 3.9
West 20.4 20.5 18.4 18.4 2.4 2.6 1.5 1.5 1.2 1.2

The measurements showing increasing power across connectors (eg after diode to before isolator on W path) are correct, and I attribute to different beam diameter launching from fiber vs from connector affecting the power meter's reading. Not including connectors, these are the expected losses of the components, with anomalous losses highlighted:

Part Expected Optical loss (without connectors, dB) Observed on E path (dB) Observed W path (dB)
IO-G-1550-APC (isolator) 0.55 0.70 0.45
MXAN-LN-10 (EOAM) 3.5 3.36 8.85
MPX-LN-0.1 (EOPM) 2.7 2.56 2.39
PN1550R2A1 (pickoff) 0.8 0.72 0.97

Obviously we should be getting a lot more power from the W path -- and it's encouraging that the laser may not be at fault! Both lasers are near their nominal output power, and in particular we probably could get no more than another 0.1 mW from the E path (assuming we could recover the extra 0.15 dB at the isolator).

I turn off the W laser to inspect the fiber tips with magnifying glasses (from EE), but nothing visibly dirty. Here is an in-depth document describing inspecting and cleaning fibers. Dry wiping the EOAM's pigtails did not change transmitted power. Perhaps we can try again with a fiber microscope for a better view. The west EOAM itself has one bent pin, and is also suspect.

aligning fiberport

Afterwards, I followed Thorlabs' steps to pre-align the W beam fiberport. The E beam fiberport appears collimated to the extent I can tell by eye, but I'd like to use the beam profiler to make fine adjustments.

quick spectrum

Since the fiber PD is set up, I measured the beat note's frequency noise with Moku's phasemeter. Saved the binary and will plot when I'm with my laptop.

exit Fri Jan 15 18:16:38 2021

ELOG V3.1.3-